# MultiModalPFN: Extending Prior-Data Fitted Networks

### v2

@@ 2 Related works
Vision–Language Multimodal Models
@@
Tabular and Multimodal Models
@@
Unstructured–structured integration has also been explored in image-centric datasets. Representative works include MMCL [23], which aligned tabular and image embeddings with contrastive learning; TIP [17], which improved robustness to incomplete features; STiL [16], which leveraged unlabeled data via semi-supervised pseudo-labeling; TIME [43], which employed TabPFN [29] as a tabular encoder; and Turbo [32], which strengthened cross-modal reasoning. Beyond individual models, toolkits such as AutoGluon [58] and modular pipelines [22] provide practical infrastructure for integrating text, image, and tabular features. Despite progress, most multimodal approaches remain concentrated on vision–language tasks, with systematic treatment of structured data still limited. Fusion strategies are often heuristic, with weak guarantees under low-data regimes or modality imbalance. Addressing these gaps is critical for building generalizable multimodal tabular systems.
General-Purpose Pre-trained Models
@@
In computer vision, self-supervised pretraining emerged as the dominant paradigm. DINOv2 [45] and DINOv3 [52] demonstrated scalable self-distillation for robust visual features, EVA [20, 19] advanced masked image modeling with large Vision Transformers, and iBOT [64] combined masking with self-distillation for effective ViT representation learning. For structured data, TabPFN [29, 30] extended this paradigm by pretraining on large synthetic datasets, learning a general prior over tabular distributions. This enables strong performance on small- and medium-sized datasets in a single forward pass without task-specific fine-tuning, positioning TabPFN as a foundation model for tabular learning. Taken together, these advances demonstrate the effectiveness of pretraining. Yet, compared to NLP and vision, pretraining for multimodal tabular data remains underexplored. Bridging this gap—particularly for multimodal integration with structured inputs—is a key step toward more comprehensive foundation models.
@@ 3 Proposed Method
Fig. 1 (a) illustrates the overall architecture of our multimodal PFN (MMPFN) that extends TabPFN [29, 30] to the multimodal setting, where image or text modalities accompany tabular inputs. Therefore, we begin by briefly reviewing TabPFN. We then describe the proposed multimodal PFN architecture, including the per-modality encoders and the modality projector that aligns non-tabular embeddings with the tabular feature space, followed by the training protocol for fine-tuning MMPFN on downstream multimodal tasks. Finally, we analyze the phenomenon of attention imbalance that arises from token-count disparities across modalities.
@@ 3.1 Preliminary: TabPFN
TabPFN [29, 30] is a tabular foundation model that treats tabular learning as amortized Bayesian inference. More specifically, a transformer is pretrained on a large collection of synthetic tabular datasets sampled from structural causal model priors. During pretraining, it learns to map a small labeled training set and an accompanying query set directly to posterior-predictive label distributions in a single forward pass. Once pretrained, TabPFN can be applied to new tabular tasks without task-specific optimization, simply by feeding the new training–test pairs to the network.
@@ 3.2 Multimodal PFN: Architecture
As shown in Fig. 1 (a), MMPFN consists of per-modality encoders, a modality projector, and a TabPFN backbone. The per-modality encoders map each input modality to a feature representation, while the modality projector aligns image and text embeddings with the shared tabular embedding space. The TabPFN backbone then jointly processes the resulting multimodal embeddings, and a lightweight decoder head produces predictions for the test samples.
Per-Modality Encoders
@@
Modality Projector
The modality projector transforms image and text embeddings into tabular-like representations, which share $d$-dimensional space compatible with the TabPFN backbone. It comprises two sublayers: a multi-head gated MLP (MGM) and a cross-attention pooler (CAP). MGM addresses the limitation of a single [CLS] embedding, which can overly compress image/text information, by expanding it into $N$ parallel $d$-dimensional projections. Fig. 1 (b) illustrates the detailed structure of MGM. Specifically, the [CLS] embedding is fed into $N$ MLP heads that project the encoder output dimension to $d$ and produce candidate modality-specific tokens. A Gated Linear Unit (GLU) [12] modulates the contribution of each head, encouraging head-wise specialization and preserving diverse aspects of the original non-tabular representation in the resulting token set.
CAP then balances tabular and non-tabular cues before fusion into the TabPFN backbone. As depicted in Figure 1 (c), it takes the $N$ MGM tokens as keys and values, and introduces $K$ learnable query vectors that attend to them via cross-attention, yielding $K$ representative $d$-dimensional embeddings per modality. The resulting pooled tokens are further refined by a MLP. These $K$ tokens form a compact and calibrated summary of the image/text information and are concatenated with the tabular tokens along the feature dimension to construct the multimodal input table for TabPFN. Without such pooling, an excessive number of non-tabular tokens can induce attention imbalance, where the modality with more tokens dominates the attention budget and suppresses tabular signal. CAP mitigates this by producing a compact, calibrated set of embeddings for the TabPFN backbone. We analyze this phenomenon in Section 3.4.
@@ 3.4 Attention Imbalance in MMPFN
Hence, when per-token quality is comparable ($c_{I}\approx c_{T}$), token-count imbalance ($N_{I}>N_{T}$) induces attention imbalance, potentially degrading performance. Consequently, MMPFN’s performance might vary with the modality token ratio. This suggests the importance of CAP. In Section 4, we empirically validate this observation by varying $K$ in the CAP.
@@
CBIS-DDSM(Calc) [50] 1546 326 8 3 5 3 0 2
Airbnb [1] 5465 1367 50 27 23 0 1 10
@@ 4.1 Experimental Setup
  • CBIS-DDSM [50] is a curated subset of the DDSM containing digitized mammography images with annotated regions of interest for calcifications and masses, biopsy-verified benign/malignant labels, and rich lesion-level metadata.
@@
  • Cloth [2] comprises 23,486 customer reviews with textual fields and 10 tabular features, providing a multimodal benchmark for sentiment and recommendation prediction.
@@
We fine-tune MMPFN for $100$ iterations to minimize the cross-entropy loss. We employ AdamW [41] with learning rate of $1\times 10^{-5}$ and batch size of $1$. For all experiments, random seeds are fixed to $\{0,1,2,3,4\}$ and then report the averaged accuracy to assess the performance of the proposed method. More implementation details are provided in the supplementary material.
Table 2: Comparison with state-of-the-arts on image-tabular multimodal datasets. Performance is reported as “accuracy (rank)”, where the accuracy are averaged over five random seeds and the rank represents the ordering of the methods according to their accuracy within each dataset (lower rank indicates better performance). “Avg.” denote the averaged accuracy across datasets. Best accuracy is shown in bold, second best in underline.
@@
TIME [43] 80.35 (5) - 72.70 (3) [^1] We cite all results of Luo et al. [43] directly. Although TIME used the CBIS-DDSM dataset without specifying subtype, the reported sample size matches the calcification subset, so we list it under CBIS-DDSM calcification in Table 2. Since the code is unavailable, reproduction was infeasible. 39.25 (2) 3.33
@@
4.2 Main Results
Results on Tabular–Image Modality Datasets.
Table 2 summarizes the classification accuracy of MMPFN and state-of-the-art baselines on four tabular–image datasets. Compared with fine-tuned TabPFN [30], which use only tabular inputs, MMPFN consistently improves performance by leveraging image features. In Table 2, MMCL [23], TIP [17], and HEALNet [28] show inconsistent performance, due to small dataset size and the low dimensionality of tabular features. In contrast, MMPFN achieves the best results on all datasets except Mass, where its performance is only marginally different from the other two models. Compared with TIME [43], which uses TabPFN as its tabular encoder, MMPFN delivers substantial gains, indicating that our modality projection strategies are more effective than simple fusion. CatBoost [14] incorporated image embeddings as raw input features, following the recommendations in its documentation, and achieved strong performance. AutoGluon [58], an AutoML framework for multimodal data, also performed competitively on several tabular benchmarks. However, both models exhibited a lower average rank than MMPFN.
@@ 4.2 Main Results
Table 3 reports results on tabular–text modality datasets. As in the image setting, adding text features consistently outperforms the fine-tuned TabPFN baseline. MMPFN is particularly strong on Airbnb. This dataset includes 50+ tabular features and a single text field, allowing tabular-specialized models to capture most of the predictive signal. Accordingly, MMPFN substantially outperforms language model–based methods, such as TFN [62] and MulT [60], which struggle to exploit the abundant tabular features. By contrast, in Cloth, the tabular part has few, weakly informative features, while the review text carries most of the signal. This is reflected in the poor performance of tabular-only models and the strong results of AllTextBERT [5], indicating that text-specialized models excel in such cases. Even so, among methods that explicitly preserve tabular structure, MMPFN achieves the best overall performance, trailing the text-specialized baseline by only a small margin. This contrasts with prior multimodal tabular studies, which emphasized tabular-dominant datasets[23]. Overall, MMPFN is not merely strong on tabular inputs. It also models unstructured modalities effectively, establishing it as a truly multimodal approach.
Table 3: Comparison with state-of-the-arts on text-tabular multimodal datasets. Performance is reported as “accuracy (rank)”, where the accuracy are averaged over five random seeds and the rank represents the ordering of the methods according to their accuracy within each dataset (lower rank indicates better performance). “Avg.” denote the averaged accuracy across datasets. Best accuracy is shown in bold, second best in underline.
@@
TabPFN [29] 46.96 (2) 44.96 (6) 55.07 (8) 36.33 (6) 5.50
Catboost [14] 43.56 (4) 40.36 (8) 59.24 (7) 35.47 (7) 6.50
AutoGluon [58] 44.60 (3) 45.24 (5) 72.07(1) 37.96 (3) 3.00
AllTextBERT [5] 30.9 (8) 44.0 (7) 68.0(2) 34.6 (8) 6.25
TFN [62] 35.7 (7) 45.8 (3) 60.1 (6) 36.8 (5) 5.25
MulT [60] 36.3 (6) 45.4 (4) 63.6 (5) 37.6 (4) 4.75
TTT [5] 38.3 (5) 47.2 (1) 65.5 (4) 38.9 (2) 3.00
MMPFN 47.78 (1) 46.17 (2) 66.26 (3) 39.04 (1) 1.75
11footnotetext: AllTextBert converts all tabular features into strings, concatenates them, and inputs the resulting sequence into DistilBERT-base-uncased for modeling, as described in [5].
Figure 2: Performance on PU20 and Cloth as a function of the number of non-tabular features. (a) Results from non-tabular–only experiments using DINOv2/ELECTRA with an MLP baseline. (b) Results from multi-modal token imbalance experiments. In both settings, the y-axis denotes accuracy, while the x-axis corresponds to the number of non-tabular input features. For (b), when applying MGM+CAP, the number of CAP heads was fixed at 24 and 4, respectively.
@@
Table 4: Ablation on the modality projector design. We compare single-head and multi-head feature extraction mechanisms (Linear/MLP, MoE, and the proposed MGM) across all datasets. MoE denotes Mixture-of-Experts. Best results are highlighted in bold.
@@
4.3 Analysis
MMPFN as an Image and Text Classifier,
Fig. 2(a) evaluates MMPFN with non-tabular–only inputs. For the baseline, we extract the pre-trained DINOv2 [45] or Electra [9] $[CLS]$ embedding and train an attached MLP head for classification—an approach that is widely adopted and often near-optimal [56]. MMPFN generated non-tabular feature embeddings from the CLS token using only MGM. With non-tabular embeddings as the sole inputs, MMPFN is within $\sim$1% of DINOv2 (69.30% vs. 69.89%). Accuracy increases consistently as the number of image feature tokens grows, indicating that additional tokens capture complementary, higher-resolution information. Thus, despite being trained with a tabular synthetic data, MMPFN functions effectively as a classifier over image embeddings cast into tabular-like features, highlighting seamless integration of non-tabular modalities with tabular data. We also note that the presence of CAP yields a modest additional gain for MMPFN.
Attention Imbalance.
We empirically investigate attention imbalance in multimodal processing. Fig. 2(a) shows the performance of MMPFN when only non-tabular features are provided as input. We use PU as the image dataset and Cloth as the text dataset. The results show that performance increases as the number of MGM heads increases, indicating that performance improves for both image and text modalities as the number of extracted non-tabular features grows. In this setting, CAP is not used to consolidate inputs. From these results, we confirm that a PFN trained solely on synthetic tabular data can classify image and text features—once projected into a tabular feature space—with performance comparable to or better than general baselines. This suggests that TabPFN has the potential to operate not only as a tabular-specific model but also as a model capable of handling multiple modalities.
However, as shown in Fig. 2(b), when tabular and non-tabular features are provided together as input, naively increasing the number of non-tabular features can degrade performance. MMPFN achieves its highest accuracy when the counts of tabular and non-tabular features are similar, and performance declines as one modality dominates the token composition. This contrasts with Fig. 2(a), where MMPFN operates in a non-tabular-only setting and does not exhibit a performance drop as the number of non-tabular features increases. This non-monotonic pattern is consistent with attention imbalance: when one modality occupies most of the input tokens, it absorbs a disproportionately large share of the attention budget. Tokens from the minority modality receive relatively little attention, diluting their useful signals, while noise features from the majority modality may still receive nonzero attention, ultimately degrading performance.
The proposed MGM+CAP mitigates this issue by first extracting non-tabular features using a sufficient number of MGM heads and then consolidating them. As a result, MGM+CAP improves steadily as the number of MGM heads increases and outperforms the MGM-only setup. The MGM head count corresponds to the x-axis, and CAP converts the extracted features into a compact, fixed-size representation—24 tokens for PU20 and 4 tokens for Cloth, preserving essential information while capping the overall token count. More detailed experimental results related to attention imbalance are provided in the supplementary material.
Modality Projector.
Table 4 summarizes an ablation of the modality projector design, comparing single-head (Linear, MLP) and multi-head (MLP, MoE, MGM) variants across all datasets. Single-head baselines, which apply a single linear or MLP projection to the non-tabular [CLS] embedding, already improve over the tabular-only backbone but yield the lowest average accuracies, highlighting the limitation of overcompressed representations.
Moving to a multi-head projector that expands [CLS] into multiple tokens consistently boosts performance, confirming the benefit of capturing diverse aspects of the image/text features. Within the multi-head family, the Mixture-of-Experts (MoE) variant is less effective and less stable across datasets, suggesting that sparse expert routing is hard to exploit in the low-data regime considered here. In contrast, the proposed MGM, which combines multi-head projection with GLU-based gating, achieves the best accuracy on every dataset and the highest overall average.
@@ 4.3 Analysis
Correlation Analysis of Cross-Modal Embeddings
Fig. 3 visualizes cosine similarities among TabPFN–backbone embeddings on all tabular–image and tabular–text datasets. These similarities illustrates the predictive relationships between features learned by MMPFN. As expected, within-modality blocks exhibit high similarity. However, several tabular–image/text pairs are also strongly aligned, indicating that MMPFN models cross-modal interactions rather than only within-modality structure. Details of the cosine similarity computation are provided in supplementary material.
Robustness in Low-Data Regimes
Table 5: Performance in Low-data Regime. Each method uses two rows: accuracy (top) and percentage change vs. full-data (bottom). The ‘Avg.’ column averages percentage changes only. Best accuracy is in bold.
@@
(-14.14) (+0.75) (-5.23) (-12.30) -10.27
Tabular datasets often require expert annotation, resulting in limited sample sizes and sparse labels [17, 16]. Consequently, models that remain robust under data scarcity are highly desirable. In such low-data regimes, MMPFN demonstrates strong performance. Table 5 compares the results of MMPFN and TIP when trained on only 10% of randomly selected samples from each dataset. While TIP employs self-supervised pretraining using all unlabeled data, our evaluation focuses on supervised finetuning under limited labeled data.
The analysis shows that, although the relative performance drop was larger for MMPFN, it consistently outperformed TIP across all datasets, even when trained with only 10% of the data. Interestingly, on the CBIS-DDSM Mass dataset, performance improved under subsampling. This behavior suggests that the PFN, pretrained on synthetic priors, can better capture discriminative characteristics when finetuned on a smaller set of labeled examples. Additional details on the behavior of MMPFN in low-data settings are provided in the supplementary material.
Scaling with Added Modalities
We assess MMPFN as multiple non-tabular modalities are added. On Petfinder, we compare against AutoGluon, a multimodal AutoML system supporting image and text modalities. As shown in Fig. 4, MMPFN’s accuracy increases monotonically from tabular $\rightarrow$ tabular+text $\rightarrow$ tabular+image $\rightarrow$ tabular+image+text (39% $\rightarrow$ 40% $\rightarrow$ 41%), indicating complementary signal from both image and text. These results have particular significance for tabular modeling, where performance improvements from architectural changes alone are often saturated. Adding complementary modalities offers a practical route to further gains. Moreover, MMPFN outperforms AutoGluon under every combination. Unlike AutoGluon’s large ensembles, MMPFN achieves higher accuracy with a lightweight and specialized architecture.

### v3

@@ 2 Related works
Vision–Language Multimodal Models.
@@
Tabular and Multimodal Models.
@@
Unstructured–structured integration has also been explored in image-centric datasets. Representative works included MMCL [20], which aligned tabular and image embeddings through contrastive learning; TIP [14], which improved robustness to missing features; STiL [13], which leveraged unlabeled data through semi-supervised pseudo-labeling; TIME [42], which used TabPFN [25] as a tabular encoder; and Turbo [28], which strengthened cross-modal reasoning. Beyond individual models, toolkits such as AutoGluon [57] and modular pipelines [19] provided practical infrastructure for multimodal integration. Despite this progress, most work remained focused on vision–language tasks, and systematic treatment of structured data remained limited. Fusion strategies were often heuristic and less reliable under low-data regimes or modality imbalance. These gaps motivated more general multimodal tabular systems.
General-Purpose Pre-trained Models.
@@
In computer vision, self-supervised pretraining has become a dominant paradigm. DINOv2 [44] and DINOv3 [51] showed scalable self-distillation for robust visual features, EVA [17, 16] advanced masked image modeling with large Vision Transformers, and iBOT [63] combined masking and self-distillation for effective ViT representations. For structured data, TabPFN [25, 26] extended this idea by pretraining on large synthetic datasets to learn a general prior over tabular distributions. It achieved strong performance on small and medium-sized datasets in a single forward pass without fine-tuning, making it a foundation model for tabular learning. Yet pretraining for multimodal tabular data remained underexplored relative to NLP and vision. Bridging this gap is important for multimodal foundation models with structured inputs.
@@ 3 Proposed Method
Figure 1 (a) illustrates the overall architecture of our multimodal PFN (MMPFN) that extends TabPFN [25, 26] to the multimodal setting, where image or text modalities accompany tabular inputs. Therefore, we begin by briefly reviewing TabPFN. We then describe the proposed multimodal PFN architecture, including the per-modality encoders and the modality projector that aligns non-tabular embeddings with the tabular feature space, followed by the training protocol for fine-tuning MMPFN on downstream multimodal tasks. Finally, we analyze the phenomenon of attention imbalance that arises from token-count disparities across modalities.
@@ 3.1 Preliminary: TabPFN
TabPFN is a tabular foundation model that treats tabular learning as amortized Bayesian inference. More specifically, a transformer is pretrained on a large collection of synthetic tabular datasets sampled from structural causal model priors. During pretraining, it learns to map a small labeled training set and an accompanying query set directly to posterior-predictive label distributions in a single forward pass. Once pretrained, TabPFN can be applied to new tabular tasks without task-specific optimization, simply by feeding the new training–test pairs to the network.
@@ 3.2 Multimodal PFN: Architecture
As shown in Figure 1 (a), MMPFN consists of per-modality encoders, a modality projector, and a TabPFN backbone. The per-modality encoders map each input modality to a feature representation, while the modality projector aligns image and text embeddings with the shared tabular embedding space. The TabPFN backbone then jointly processes the resulting multimodal embeddings, and a lightweight decoder head produces predictions for the test samples.
Per-Modality Encoders.
@@
Modality Projector.
The modality projector transforms image and text embeddings into tabular-like representations, which share $d$-dimensional space compatible with the TabPFN backbone. It comprises two sublayers: a multi-head gated MLP (MGM) and a cross-attention pooler (CAP). MGM addresses the limitation of a single [CLS] embedding, which can overly compress image/text information, by expanding it into NN parallel $d$-dimensional projections. Figure 1 (b) illustrates the detailed structure of MGM. Specifically, the [CLS] embedding is fed into NN MLP heads that project the encoder output dimension to $d$ and produce candidate modality-specific tokens. A Gated Linear Unit (GLU) [10] modulates the contribution of each head, encouraging head-wise specialization and preserving diverse aspects of the original non-tabular representation in the resulting token set.
CAP then balances tabular and non-tabular cues before fusion in the TabPFN backbone. As shown in Figure 1 (c), it takes the NN MGM tokens as keys and values and introduces $K$ learnable query vectors that cross-attend to them, yielding $K$ representative $d$-dimensional embeddings per modality. The pooled tokens are refined by an MLP. These $K$ tokens form a compact, calibrated summary of image/text information and are concatenated with the tabular tokens along the feature dimension to construct the multimodal input table for TabPFN. Without pooling, too many non-tabular tokens can induce attention imbalance, where the modality with more tokens dominates the attention budget and suppresses tabular signal. CAP mitigates this by producing a compact, calibrated set of embeddings for the TabPFN backbone. We analyze this in Section 3.4.
@@ 3.4 Attention Imbalance in MMPFN
Hence, when per-token quality is comparable ($c_{I}\approx c_{T}$), token-count imbalance ($N_{I}>N_{T}$) induces attention imbalance, potentially degrading performance. Consequently, MMPFN’s performance might vary with the modality token ratio. This suggests the importance of CAP. In Section 4, we validate this observation by varying $K$ in the CAP.
@@
CBIS-DDSM(Calc) [49] 1545 326 8 3 5 3 0 2
Airbnb [30] 18316 4579 50 27 23 0 1 10
@@ 4.1 Experimental Setup
  • CBIS-DDSM [49] is a curated subset of DDSM with digitized mammograms, annotated regions of interest for calcifications and masses, biopsy-verified benign/malignant labels, and lesion-level metadata.
@@
  • Cloth [32] comprises 23,486 customer reviews with text fields and 10 tabular features for multimodal sentiment and recommendation prediction.
@@
We fine-tune MMPFN for $100$ iterations using cross-entropy loss. We use AdamW [40] with a learning rate of $1\times 10^{-5}$ and batch size 11. For all experiments, we use random seeds $\{0,1,2,3,4\}$ and report average accuracy. More details are provided in the supplementary material.
00footnotetext: We cite all results of Luo et al. [42] directly. Although TIME used the CBIS-DDSM dataset without specifying subtype, the reported sample size matches the calcification subset, so we list it under CBIS-DDSM calcification in Table 2. Since the code is unavailable, reproduction was infeasible.
4.2 Main Results
Results on Tabular–Image Modality Datasets.
Table 2 summarizes the classification accuracy of MMPFN and state-of-the-art baselines on four tabular–image datasets. Compared with fine-tuned TabPFN [26], which uses only tabular inputs, MMPFN consistently improves performance by leveraging image features. MMCL [20], TIP [14], and HEALNet [24] show inconsistent results, likely due to the small dataset size and low-dimensional tabular features. In contrast, MMPFN achieves the best results on all datasets except Mass, where it remains competitive with the top models. Compared with TIME [42], which uses TabPFN as its tabular encoder, MMPFN delivers substantial gains, suggesting that our modality projection strategies are more effective than simple fusion. CatBoost [47] uses image embeddings as raw input features and achieves strong performance. AutoGluon [57], an AutoML framework for multimodal data, also performs competitively on several benchmarks. However, MMPFN achieves a better average rank than both models.
Table 2: Comparison with state-of-the-art on tabular-image multimodal datasets. Results are reported as accuracy (rank), averaged over five random seeds, where lower rank indicates better performance. “Avg.” denotes the mean accuracy across datasets. Best and second-best results are marked in bold and underline.
@@ 4.2 Main Results
TIME [42] 80.35 (5) - 72.70 (3) [^1] AllTextBert converts all tabular features into strings, concatenates them, and inputs the resulting sequence into DistilBERT-base-uncased for modeling, as described in [3]. 39.25 (2) 3.33
@@
Table 3 reports results on tabular–text datasets. As in the image setting, adding text features consistently improves over the fine-tuned TabPFN baseline. MMPFN is particularly strong on Airbnb, which includes 50+ tabular features and a single text field, allowing tabular-specialized models to capture most of the predictive signal. Accordingly, MMPFN substantially outperforms language model–based methods, such as TFN [61] and MulT [59], which struggle to exploit abundant tabular features. By contrast, Cloth has few informative tabular features, while the review text carries most of the signal. This is reflected in the weak performance of tabular-only models and the strong results of AllTextBERT [3], indicating that text-specialized models excel in such cases. Even so, among methods that explicitly preserve tabular structure, MMPFN achieves the best overall performance, trailing the text-specialized baseline by only a small margin. This contrasts with prior multimodal tabular studies, which focused on tabular-dominant datasets [20]. Overall, MMPFN effectively handles both tabular and unstructured modalities.
Table 3: Comparison with state-of-the-art on tabular-text multimodal datasets. Results are reported as accuracy (rank), averaged over five random seeds, where lower rank indicates better performance. “Avg.” denotes the mean accuracy across datasets. Best and second-best results are marked in bold and underline.
@@
TabPFN [25] 46.96 (2) 44.96 (6) 55.07 (9) 36.33 (7) 6.00
Catboost [47] 43.56 (4) 40.36 (9) 59.24 (8) 35.47 (8) 7.25
AutoGluon [57] 44.60 (3) 45.24 (5) 72.07 (1) 37.96 (4) 3.25
AllTextBERT [3] 30.9 (9) 44.0 (7) 68.0 (3) 34.6 (9) 7.00
TFN [61] 35.7 (8) 45.8 (3) 60.1 (7) 36.8 (6) 6.00
MulT [59] 36.3 (7) 45.4 (4) 63.6 (6) 37.6 (5) 5.50
TTT [3] 38.3 (6) 47.2 (1) 65.5 (5) 38.9 (3) 3.75
TabSTAR [1] 40.06 (5) 43.75 (8) 71.75 (2) 41.53 (1) 4.00
MMPFN 47.78 (1) 46.17 (2) 66.26 (4) 39.04 (2) 2.25
4.3 Analysis
MMPFN as an Image and Text Classifier.
Figure 2 (a) evaluates MMPFN with non-tabular–only inputs. As a baseline, we use pre-trained DINOv2 [44] or Electra [7] $[CLS]$ embeddings with an attached MLP classifier, a widely used and often near-optimal setup [55]. MMPFN uses only MGM to generate non-tabular embeddings from the CLS token. With these embeddings as the sole inputs, MMPFN remains within $\sim$1% of DINOv2 (69.30% vs. 69.89%). Accuracy increases with the number of image tokens, suggesting that additional tokens capture complementary, higher-resolution information. Although trained on synthetic tabular data, MMPFN effectively classifies image embeddings mapped to tabular-like features, showing that it is not limited to native tabular inputs. CAP further provides a modest gain.
Attention Imbalance.
We empirically analyze attention imbalance in multimodal processing. In Figure 2 (a), MMPFN uses only non-tabular inputs. Accuracy increases with the number of MGM heads, showing that additional non-tabular tokens improve performance. CAP is not used here. These results show that a PFN trained on synthetic tabular data can classify projected non-tabular features competitively with standard baselines. TabPFN is therefore not limited to native tabular inputs.
Figure 2: Performance on PU20 versus the number of non-tabular tokens. (a) Image-only results with DINOv2 and an MLP baseline. (b) Multimodal results under token imbalance. The y-axis shows accuracy and the x-axis shows the number of MGM heads. In (b), MGM+CAP uses 24 CAP heads.
Figure 3: Token count and attention mass. Attention mass for tabular and non-tabular tokens is measured as the number of non-tabular tokens varies, without CAP. Values are averaged over 12 self-attention layers in TabPFN. PU20 and Salary use 11 and 4 tabular tokens. The x-axis is in log scale.
In Figure 2 (b), the trend changes when tabular and non-tabular features are used together. MMPFN performs best when the two modalities have similar token counts, and performance drops as one modality dominates the sequence. This pattern does not appear in Figure 2 (a), where only non-tabular inputs are used. The contrast is consistent with attention imbalance: the modality with more tokens absorbs more of the attention budget, while the other receives less attention and contributes less signal. MGM+CAP mitigates this problem by extracting non-tabular features with enough MGM heads and then compressing them into 24 CAP tokens. As the number of MGM heads increases, MGM+CAP improves steadily and outperforms MGM alone.
We further examine attention allocation by varying the number of non-tabular tokens while keeping the number of tabular tokens fixed. In Figure 3, attention mass shifts monotonically toward the non-tabular modality as its token count increases, while the mass on tabular tokens decreases. This result supports the same explanation: a modality with more tokens receives a larger share of the attention budget. Table 4 then separates the effects of token count and representation quality. Increasing token count alone reduces MGM by nearly 2%p despite identical representations, whereas MGM+CAP remains stable. When representation quality improves at the same total token count ($4\times 32$ and 128), MGM still suffers from the larger token set, but MGM+CAP benefits from the stronger representations by compressing them into a compact set of tokens. More details are provided in the supplementary material.
[…]
MGM 84.17 82.43 83.91
MGM+CAP 84.10 83.57 84.59
[…]
MGM 63.12 61.76 61.86
MGM+CAP 64.20 64.22 65.03
Modality Projector.
Table 5 ablates the modality projector, comparing single-head (Linear, MLP) and multi-head (MLP, MoE, MGM) variants. Parameter budgets are provided in the supplementary material. Single-head baselines apply one linear or MLP projection to the non-tabular [CLS] embedding. Although they improve over the tabular-only backbone, they yield the lowest average accuracies, suggesting that a single projection overcompresses non-tabular information. Expanding [CLS] into multiple tokens consistently improves performance, showing the benefit of capturing diverse aspects of image/text features. Within this family, MoE is less effective and less stable across datasets, suggesting that sparse expert routing is hard to exploit in the low-data regime. By contrast, MGM combines multi-head projection with GLU-based gating and achieves the best accuracy on every dataset and the best overall average.
We next examine how the modality projector incorporates non-tabular information. Specifically, we compare CAP, which pools non-tabular features into representative tokens, with Feature-wise Linear Modulation (FiLM) [46], which uses non-tabular representations to generate feature-wise affine parameters for tabular tokens. As shown in Table 6, CAP consistently outperforms FiLM on all datasets. This result suggests that controlling token count is important for tabular–non-tabular fusion because it alleviates attention imbalance between the two modalities. FiLM applies the same channel-wise transformation to all tabular tokens, limiting token-specific use of non-tabular cues. CAP instead compresses non-tabular features into representative tokens that the TabPFN encoder attends jointly with tabular tokens, preserving relevant cross-modal cues while reducing token imbalance.
@@ 4.3 Analysis
Table 5: Ablation of the design of MGM. We compare single-head and multi-head feature extraction mechanisms (Linear/MLP, MoE, and the proposed MGM) across all datasets. MoE denotes Mixture-of-Experts. Best results are highlighted in bold.
@@
Table 6: Ablation of the design of CAP. We compare two mechanisms for incorporating non-tabular information (FiLM and the proposed CAP) on all datasets. FiLM denotes Feature-wise Linear Modulation. Best results are highlighted in bold.
Method PU20 Mass Calc Cloth Salary Airbnb PetFinder-I PetFinder-T PetFinder-A Avg.
FiLM [46] 80.00 73.02 73.25 65.82 42.27 46.06 39.84 38.31 40.77 55.48
CAP 85.22 74.53 75.40 66.26 46.17 47.78 40.70 39.04 41.19 57.37
@@
Cross-Modal Correlation.
Figure 4 visualizes cosine similarities among TabPFN–backbone embeddings on all tabular–image and tabular–text datasets. These similarities illustrates the predictive relationships between features learned by MMPFN. As expected, within-modality blocks exhibit high similarity. However, several tabular–image/text pairs are also strongly aligned, indicating that MMPFN models cross-modal interactions rather than only within-modality structure. Details of the cosine similarity computation are provided in supplementary material.
Table 7: Performance in Low-data Regime. Each method uses two rows: accuracy (top) and percentage change vs. full-data (bottom). The ‘Avg.’ column averages percentage changes only. Best results are highlighted in bold.
@@
(-14.14) (+0.75) (-5.23) (-12.30) (-10.27)
Robustness in Low-Data Regimes.
Tabular datasets often require expert annotation, leading to limited sample sizes and sparse labels [14, 13]. Models that remain robust under such data scarcity are therefore desirable. In this setting, MMPFN performs strongly. Table 7 compares MMPFN and TIP when trained on only 10% of randomly selected samples from each dataset. Although TIP uses self-supervised pretraining on all unlabeled data, we focus on supervised finetuning with limited labeled data.
Although MMPFN shows a larger relative drop, it consistently outperforms TIP on all datasets, even with only 10% of the data. On CBIS-DDSM Mass, performance even improves under subsampling. This result suggests that the PFN, pretrained on synthetic priors, can better capture discriminative characteristics when finetuned on fewer labeled examples. More details on low-data behavior are provided in the supplementary material.
Scaling with Added Modalities.
We assess MMPFN as multiple non-tabular modalities are added. On Petfinder, we compare against AutoGluon, a multimodal AutoML system supporting image and text modalities. As shown in Figure 5, MMPFN’s accuracy increases monotonically from tabular $\rightarrow$ tabular+text $\rightarrow$ tabular+image $\rightarrow$ tabular+image+text (39% $\rightarrow$ 40% $\rightarrow$ 41%), indicating complementary signal from both image and text. These results have particular significance for tabular modeling, where performance improvements from architectural changes alone are often saturated. Adding complementary modalities offers a practical route to further gains. Moreover, MMPFN outperforms AutoGluon under every combination. Unlike AutoGluon’s large ensembles, MMPFN achieves higher accuracy with a lightweight and specialized architecture.